Financial data for Joel de Paris, Inc., for last year follow:

Joel de Paris, Inc.
Balance Sheet
Beginning
Balance
Ending
Balance
Assets
Cash $ 139,000 $ 138,000
Accounts receivable 348,000 479,000
Inventory 579,000 482,000
Plant and equipment, net 784,000 791,000
Investment in Buisson, S.A. 405,000 434,000
Land (undeveloped) 251,000 245,000
Total assets $ 2,506,000 $ 2,569,000
Liabilities and Stockholders' Equity
Accounts payable $ 373,000 $ 350,000
Long-term debt 1,008,000 1,008,000
Stockholders' equity 1,125,000 1,211,000
Total liabilities and stockholders' equity $ 2,506,000 $ 2,569,000


Joel de Paris, Inc.
Income Statement
Sales $ 4,114,000
Operating expenses 3,496,900
Net operating income 617,100
Interest and taxes:
Interest expense $ 117,000
Tax expense 191,000 308,000
Net income $ 309,100


The company paid dividends of $223,100 last year. The “Investment in Buisson, S.A.,” on the balance sheet represents an investment in the stock of another company. The company's minimum required rate of return of 15%.

Required:

1. Compute the company's average operating assets for last year.

2. Compute the company’s margin, turnover, and return on investment (ROI) for last year. (Do not round intermediate calculations and round your final answers to 2 decimal places.)

3. What was the company’s residual income last year?
Answer #1
Answer

1

Beginning operating assets = 2506000-405000-251000= $     1,850,000
Ending operating assets = 2569000-434000-252000= $     1,883,000
Average Operating Assets =(1850000+1883000)/2 $     1,866,500
2
Margin =Net operating income/Sales = 617100/4114000= 15%
Turnover = Sales/Average operating assets = 4114000/1866500= 2.20
ROI = Margin*Turnover = 15%*2.2= 33%
3
Residual income = 617100-(1866500*15%)= $ 337,525